# The impact of $\gamma N$ and $\gamma^{*}N$ interactions on our understanding of nucleon excitations

### v1

@@ 4   Baryons and QCD
In Sec. 3 we discussed the experimentally known baryon spectrum. As we have seen, the spectrum carries clear traces of the nonrelativistic quark model, which was historically also one of the pillars for the development of QCD. In quark models, the light $u$ and $d$ quarks have masses of about $300\dots 350$ MeV, which are used to reproduce the masses of ground-state baryons and their excitations. These massive quarks are called constituent quarks. On the other hand, the light current quarks in the QCD Lagrangian have masses of a few MeV, which also appear in low-energy relations of QCD [597] such as the Gell-Mann-Oakes-Renner relation [598].
The mass gap between current and constituent quarks can be explained by the spontaneous breaking of chiral symmetry [599, 600]. This mechanism dynamically generates a constituent mass through the interactions of quarks and gluons. It adds roughly $\sim 350$ MeV to the quark mass of each quark flavor, which is a subleading or even negligible effect for heavy charm and bottom quarks while for light $u$ and $d$ quarks it is the dominant contribution. An important consequence is the large mass gap between chiral partners: The masses of the nucleon with spin-parity $J^{P}=1/2^{+}$ and its chiral partner $N(1535)\nicefrac{{1}}{{2}}^{-}$ with $J^{P}=1/2^{-}$ differ by about 600 MeV. Since quark models provide no insight how this large mass difference is generated, we should take a step back and ask: How much can we infer about the baryon spectrum from QCD?
@@ 4.1   What is a baryon in QCD?
    $V_{a}^{\mu}=\hskip 1.0pt\overline{\psi}\hskip 1.0pt\gamma^{\mu}\mathsf{t}_{a}\psi\,,\qquad\partial_{\mu}V_{a}^{\mu}=i\hskip 1.0pt\overline{\psi}\hskip 1.0pt\left[\mathsf{M},\mathsf{t}_{a}\right]\psi\,,\qquad Q_{a}^{V}(t)=\int d^{3}x\,\psi^{\dagger}\mathsf{t}_{a}\psi\,.$ (4.3)
@@
    $\tilde{A}_{a}^{\mu}=\hskip 1.0pt\overline{\psi}\hskip 1.0pt\gamma^{\mu}\gamma_{5}\mathsf{t}_{a}\psi\,,\qquad\partial_{\mu}\tilde{A}_{a}^{\mu}=i\hskip 1.0pt\overline{\psi}\hskip 1.0pt\left\{\mathsf{M},\mathsf{t}_{a}\right\}\gamma_{5}\psi$ (4.4)
@@
$Z=\int\mathcal{D}A\,\mathcal{D}\psi\,\mathcal{D}\hskip 1.0pt\overline{\psi}\hskip 1.0pt\,e^{iS}\,.$ (4.5)
Here, the path integral measure $\mathcal{D}A\,\mathcal{D}\psi\,\mathcal{D}\hskip 1.0pt\overline{\psi}\hskip 1.0pt$ amounts to a statistical average over all possible quark and gluon field configurations. The central quantities in QCD are the $n$-point correlation functions of the form
@@
where $\phi\in\{A^{\mu},\psi,\hskip 1.0pt\overline{\psi}\hskip 1.0pt\}$ is a shorthand for the quark and gluon fields and $\mathsf{T}$ denotes time ordering. Abbreviating the combinations of fields by $\mathcal{O}=\phi(x_{1})\dots\phi(x_{n})$ and the respective correlation function by $\langle\mathcal{O}\rangle=G(x_{1},\dots,x_{n})$, they can be computed from the path integral via
$\langle\mathcal{O}\rangle=\frac{1}{Z}\int\mathcal{D}A\,\mathcal{D}\psi\,\mathcal{D}\hskip 1.0pt\overline{\psi}\hskip 1.0pt\,e^{iS}\,\mathcal{O}\,.$ (4.7)
@@
Fig. 4.2 shows some of the simplest correlation functions: the elementary quark propagator from the Lagrangian turns into the full (‘dressed’) quark propagator $S_{\alpha\beta}(x)=\langle\hskip 1.0pt\overline{\psi}\hskip 1.0pt_{\alpha}(x)\,\psi_{\beta}(0)\rangle$, the elementary gluon propagator turns into the dressed gluon propagator $D^{\mu\nu}(x)=\langle A^{\mu}(x)A^{\nu}(0)\rangle$, the vertices become dressed vertices and so on. In addition, there are new correlation functions which do not have an elementary counterpart, such as a quark-antiquark ($q\bar{q}$) four-point function, a three-quark correlation function, etc. Moreover, the fields in Eq. (4.6) do not need to be elementary fields but can also be composite: For example, by contraction with appropriate Dirac-flavor-color matrices, the three-quark correlator turns into a nucleon current correlator $\langle N(x)\,\bar{N}(0)\rangle$. This is especially useful for constructing gauge-invariant correlation functions: While the correlation functions with open quark or gluon legs depend on the gauge, hadronic current correlators are by construction gauge-invariant and thus in principle measurable.
@@
An important relation in QFT says that physical states produce poles in correlation functions (see e.g. [601] for a proof). Suppose we split the $n$ coordinates in Eq. (4.6) into two sets $\{x_{i}\}$ and $\{x_{j}\}$ with $i=1\dots N$. Then, by inserting the completeness relation of the state space, each asymptotic one-particle state $|\lambda\rangle$ with onshell momentum $P$ and mass $m_{\lambda}$ (such that $P^{2}=m_{\lambda}^{2}$) produces a pole in any correlation function that is compatible with the quantum numbers of the state:
@@
which is the overlap of the vacuum and the one-particle state. Here, an overall coordinate $z$ has been factored out by translation invariance, so the BSWF depends only on $N-1$ coordinates. $\Psi$ is the QFT analogue of a quantum-mechanical wave function since it carries the information about the hadron, although it does not have a direct probability interpretation (more on that in Sec. 4.5).
@@
Figure 4.3: Left: Poles and cuts in the complex energy plane of a correlation function. By tuning certain control parameters, resonances may turn into bound states and vice versa. Right: Baryon poles appear in any correlation function that can produce them, such as the $N\pi$ scattering amplitude, the gauge-dependent $qqq$ correlation function or the respective gauge-invariant current correlator.
Is some given baryon resonance more like a three-quark state or a meson-baryon dominated state? A priori, QFT does not directly tell us that: From Fig. 4.3, a nucleon resonance can be thought of as an $N\pi$ state, but also as a three-quark state ($qqq$), or a five-quark state ($qqqq\bar{q}$) etc., because all those correlation function must produce the same spectrum. For example, any dynamically generated ‘additional’ baryon resonance that arises from meson-baryon interactions in an EFT must also have a three-quark core (unless some quantum number prevents it), so these two options are not binary. Vice versa, a pole in the $qqq$ correlation function emerging from quark-gluon interactions must also appear in the $qqqq\bar{q}$ and other correlation functions, but the three-quark component might be small.
The relevant question is therefore: How much overlap does a state have with a given configuration? This is defined by the respective BSWF. If the $qqq$ overlap is small and the $qqqq\bar{q}$ overlap large, it implies a dominant $qqqq\bar{q}$ (or $N\pi$) structure; if the $qqq$ overlap is large it implies a three-quark dominance. In light-quark baryon spectroscopy, the $\Lambda(1405)$ is the primary candidate for having a dynamically generated additional state, which would then translate into two poles that couple more strongly to $qqq$ or $qqqq\bar{q}$ [602, 603, 604, 605]. Going further, a deeper question is: What is the reason for the existence of a particular resonance? What are the interactions that produce a $qqq$-like or molecular behavior? Is there a molecular force that binds $N$ and $\pi$, and this bound state may also couple to $qqq$? Or is there a $qqq$ bound state that acquires a molecular component? Microscopic calculations that can extract such information using different approaches are underway and will presumably shed more light on these questions in the future. In addition, much can be learned from the structure of baryons as measured in electroproduction, which we will discuss in Sec. 5.
@@ 4.2   Effective field theories
Effective field theories (EFTs) for describing hadron-hadron interactions have found widespread applications over the last decades, see e.g. [606, 607, 604, 460] for recent reviews. The primary quantities of interest are hadronic scattering amplitudes like in the top right row of Fig. 4.3. These contain the hadron poles matching the quantum numbers allowed by the initial and final states, which allows one to extract the hadron spectrum from the pole positions. Generally, the analytic structure of scattering amplitudes is constrained by the S-matrix principles of unitarity, analyticity and crossing symmetry. The analytic structure is determined entirely by physical particles: bound states and resonances manifest as poles, while particles in the loop diagrams generate branch cuts. By contrast, confinement implies that quarks and gluons must conspire in such a way that all direct evidence of their existence is erased from the singularity structure of physical scattering amplitudes. Although conceptually inevitable, the elimination of these gauge-dependent degrees of freedom from gauge-invariant amplitudes is quite astonishing and a central theme of quark-hadron duality [608].
@@
In practice the situation can be much more complicated than this simple sketch. For example, the case with two branch points is shown in Fig. 4.5. The four Riemann sheets in the left panel can be mapped to a torus [604] in the right panel: The bottom and top borders of the rectangle are identified, and the left and right borders correspond to $|s|=\infty$. Crossing over between the four sheets is now an analytic operation.
@@
$\tau(s)=\frac{1}{16\pi}\sqrt{\frac{s-4m^{2}}{s}}=\frac{1}{16\pi}\frac{2|{\mbox{\boldmath$q$}}|}{\sqrt{s}}\,,$ (4.12)
and $q$ is the three-momentum in the CM frame (see also [609] for a pedagogical introduction). Unitarity thus constrains the imaginary parts of the partial waves $f_{l}(s)$. This can be expressed through the phase shift $\delta_{l}(s)$: Writing $f_{l}(s)=R_{l}(s)\,e^{i\delta_{l}(s)}$ above the threshold $\sqrt{s}>2m$, Eq. (4.11) implies $R_{l}=\frac{1}{\tau}\sin\delta_{l}$ and hence
@@
It is also often useful to perform a split $\mathbf{G}_{0}=\mathbf{G}_{1}+\mathbf{G}_{2}$ and define a quantity $\mathbf{K}_{1}^{-1}=\mathbf{K}^{-1}-\mathbf{G}_{1}$Leftrightarrow $\mathbf{K}_{1}=\mathbf{K}+\mathbf{K}\,\mathbf{G}_{1}\,\mathbf{K}_{1}$, so that Eq. (4.15) turns into
@@
How is the kernel $\mathbf{K}$ determined, which contains the interactions between the particles and is the input in the equations? Since the particles appearing in the loops in Fig. (4.6) are hadrons, these interactions cannot be the ones between quarks and gluons but should rather be derived from an effective Lagrangian containing physical fields like nucleons, pions, etc. This leads to EFTs, whose prime example is chiral perturbation theory (ChPT) [610, 611, 597]. In the two-flavor case, the ChPT Lagrangian is expressed in terms of pions which couple to nucleons; the extension to three flavors is straightforward. The idea behind ChPT is to be agnostic about the microscopic interactions and include all possible terms compatible with the symmetries of QCD – in particular chiral symmetry – which leads to a systematic expansion in powers of derivatives and pion masses. The elementary Feynman diagrams are shown in Fig. 4.7 together with the full correlation functions like the nucleon and pion propagators, $N\pi$ and $\pi\pi$ scattering amplitudes, etc., which can be calculated through perturbative loop expansions. The price to pay is that the Lagrangian is non-renormalizable and contains infinitely many terms with infinitely many free parameters. However, the derivative couplings ensure that higher loop diagrams in the perturbative expansion are suppressed at low momenta and small pion masses. The renormalization can then be done at each order and requires only a finite number of parameters at each order, the low-energy constants, which must be determined from elsewhere, e.g., from experiment or lattice QCD. In this way, ChPT is a low-energy EFT and one can make a range of predictions at low momenta and small pion masses, e.g. for the nucleon mass as a function of the pion mass, the $N\pi$ scattering amplitude, electroweak processes and so on, see [612, 613, 614, 615, 616, 617, 618, 619] for reviews.
@@
On the other hand, because ChPT amounts to perturbative loop expansions it cannot describe bound states and resonances. One strategy to deal with this is to include these extra fields ($\rho$ and $\omega$ mesons, the $\Delta$ resonance, Roper resonance, $\Lambda(1405)$, etc.) explicitly in the Lagrangian and hence in the perturbative expansions, see e.g. [620, 621, 622, 623, 624, 625, 626, 627, 628, 629, 630] and references therein. Another strategy is to unitarize ChPT by employing BSEs [631, 632, 633, 634, 635, 636]. A BSE automatically produces the unitary cuts when the hadrons in the loops go onshell, but because it is genuinely non-perturbative it can also dynamically generate bound states and resonances. This non-perturbative nature can be understood by analogy with the geometric series: The series expansion $f(x)=1+x+x^{2}+x^{3}+\dots$ determines the function $f(x)=1/(1-x)$ only for $|x|<1$, whereas the ‘nonperturbative’ equation
@@
where $f(x)$ also appears on the right, is always exact irrespective of how large $x$ is. Similarly, for small couplings one can expand Eq. (4.15) in $\mathbf{T}=\mathbf{K}+\mathbf{K}\,\mathbf{G}_{0}\,\mathbf{K}+\dots$, which does not generate any poles, while the original nonperturbative equation generates poles if the condition ‘$\mathbf{K}\,\mathbf{G}_{0}=1$’ is satisfied for some value of $\sqrt{s}=m_{\lambda}$.
@@
A shortcoming of the BSE in the form of Eq. (4.15), which is only iterated in the $s$ channel, is that it does not automatically satisfy crossing symmetry. While it generates the correct $s$-channel cut for $s>4m^{2}$ (the unitary or right-hand cut), a unitarization also in the crossed channels would produce further cuts in the Mandelstam variables $t$ and $u$, which translates to $s<0$ and is the so-called left-hand cut. For this reason, other unitarization methods based on dispersion relations such as the $N/D$ approach [637, 638, 639], the inverse amplitude method [640, 641, 642, 643] or the Roy-Steiner equations [644, 645, 646, 647] have been frequently employed, see e.g. the reviews [643, 644, 604] for details and references.
@@
The application of BSEs with kernels taken from EFTs, with (multi-)pion and possibly other hadron exchanges, has found widespread applications in hadron and nuclear physics. Although we exemplified it for the two-body scattering amplitude in a single channel, the approach can be generalized to dynamical coupled-channel equations where the external particles create different intermediate states such as $N\pi$, $\Lambda K$, $\Sigma K$, $N\gamma$, etc., see [460] for a recent review. Coupled-channel models have become indispensable tools for amplitude analyses to extract experimental data, especially in view of extracting the photo- and electrocouplings of light baryons (see also the discussion in Sec. 2.2.1). Examples are the ANL-Osaka approach [429, 433, 432, 437, 442], the Jülich-Bonn-Washington approach [453, 454, 455, 648, 446, 649, 460], and various K-matrix formulations like the MAID [490, 382, 376], SAID [492, 368], JLab [15], KSU [41, 650], Giessen [396, 397, 401, 402] and Bonn-Gatchina approaches [409, 414, 419, 283, 475].
@@
Dynamical coupled-channel models also allow one to estimate the meson-baryon dressing effects of nucleon resonances and extract their pole trajectories in the complex plane, see e.g. [437, 605]. As an example, the procedure employed in the ANL-Osaka approach is illustrated in Fig. 4.8 for a single-channel $N\pi$ scattering equation. If one pulls out the resonant $s$-channel $N^{\ast}$ contributions from the kernel explicitly, then also the T-matrix turns into the sum of a non-resonant and a resonant part. The former satisfies the scattering equation with the non-resonant kernel, while the dressing effects for the resonance coupling and propagator are calculated separately. In this way one can study the effects of meson-baryon interactions: When the full T-matrix is fitted to the electroproduction data, switching off the dressing effects for the masses and transition form factors provides an estimate for the ‘quark core’ of the resonance.

### v4

@@ 4  Baryons and QCD
In Sec. 3 we discussed the experimentally known baryon spectrum. As we have seen, the spectrum carries clear traces of the nonrelativistic quark model, which was historically also one of the pillars for the development of QCD. In quark models, the light $u$ and $d$ quarks have masses of about $300\dots 350$ MeV, which are used to reproduce the masses of ground-state baryons and their excitations. These massive quarks are called constituent quarks. On the other hand, the light current quarks in the QCD Lagrangian have masses of a few MeV, which also appear in low-energy relations of QCD Weinberg:1978kz such as the Gell-Mann-Oakes-Renner relation Gell-Mann:1968hlm .
The mass gap between current quarks at large momenta and constituent quarks at low momenta can be explained by the spontaneous breaking of chiral symmetry Nambu:1960tm , Goldstone:1961eq . This mechanism dynamically generates a constituent mass through the interactions of quarks and gluons. It adds roughly $\sim 350$ MeV to the quark mass of each quark flavor, which is a subleading or even negligible effect for heavy charm and bottom quarks while for light $u$ and $d$ quarks it is the dominant contribution. An important consequence is the large mass gap between chiral partners: The masses of the nucleon with spin-parity $J^{P}=1/2^{+}$ and its chiral partner $N(1535)\nicefrac{{1}}{{2}}^{-}$ with $J^{P}=1/2^{-}$ differ by about 600 MeV. Since quark models provide no insight how this large mass difference is generated, we should take a step back and ask: How much can we infer about the baryon spectrum from QCD?
@@ 4.1  What is a baryon in QCD?
    $V_{a}^{\mu}=\hskip 1.0pt\overline{\hskip-1.0pt\psi\hskip-1.0pt}\hskip 1.0pt\gamma^{\mu}\mathsf{t}_{a}\psi\,,\qquad\partial_{\mu}V_{a}^{\mu}=i\hskip 1.0pt\overline{\hskip-1.0pt\psi\hskip-1.0pt}\hskip 1.0pt\left[\mathsf{M},\mathsf{t}_{a}\right]\psi\,,\qquad Q_{a}^{V}(t)=\int d^{3}x\,\psi^{\dagger}\mathsf{t}_{a}\psi\,.$ (4.3)
@@
    $\tilde{A}_{a}^{\mu}=\hskip 1.0pt\overline{\hskip-1.0pt\psi\hskip-1.0pt}\hskip 1.0pt\gamma^{\mu}\gamma_{5}\mathsf{t}_{a}\psi\,,\qquad\partial_{\mu}\tilde{A}_{a}^{\mu}=i\hskip 1.0pt\overline{\hskip-1.0pt\psi\hskip-1.0pt}\hskip 1.0pt\left\{\mathsf{M},\mathsf{t}_{a}\right\}\gamma_{5}\psi$ (4.4)
@@
$Z=\int\mathcal{D}A\,\mathcal{D}\psi\,\mathcal{D}\hskip 1.0pt\overline{\hskip-1.0pt\psi\hskip-1.0pt}\hskip 1.0pt\,e^{iS}\,.$ (4.5)
Here, the path integral measure $\mathcal{D}A\,\mathcal{D}\psi\,\mathcal{D}\hskip 1.0pt\overline{\hskip-1.0pt\psi\hskip-1.0pt}\hskip 1.0pt$ amounts to a statistical average over all possible quark and gluon field configurations. The central quantities in QCD are the $n$-point correlation functions of the form
@@
where $\phi\in\{A^{\mu},\psi,\hskip 1.0pt\overline{\hskip-1.0pt\psi\hskip-1.0pt}\hskip 1.0pt\}$ is a shorthand for the quark and gluon fields and $\mathsf{T}$ denotes time ordering. Abbreviating the combinations of fields by $\mathcal{O}=\phi(x_{1})\dots\phi(x_{n})$ and the respective correlation function by $\langle\mathcal{O}\rangle=G(x_{1},\dots,x_{n})$, they can be computed from the path integral via
$\langle\mathcal{O}\rangle=\frac{1}{Z}\int\mathcal{D}A\,\mathcal{D}\psi\,\mathcal{D}\hskip 1.0pt\overline{\hskip-1.0pt\psi\hskip-1.0pt}\hskip 1.0pt\,e^{iS}\,\mathcal{O}\,.$ (4.7)
@@
Fig. 4.2 shows some of the simplest correlation functions: the elementary quark propagator from the Lagrangian turns into the full (‘dressed’) quark propagator $S_{\alpha\beta}(x)=\langle\hskip 1.0pt\overline{\hskip-1.0pt\psi\hskip-1.0pt}\hskip 1.0pt_{\alpha}(x)\,\psi_{\beta}(0)\rangle$, the elementary gluon propagator turns into the dressed gluon propagator $D^{\mu\nu}(x)=\langle A^{\mu}(x)A^{\nu}(0)\rangle$, the vertices become dressed vertices and so on. In addition, there are new correlation functions which do not have an elementary counterpart, such as a quark-antiquark ($q\bar{q}$) four-point function, a three-quark correlation function, etc. Moreover, the fields in Eq. (4.6) do not need to be elementary fields but can also be composite: For example, by contraction with appropriate Dirac-flavor-color matrices, the three-quark correlator turns into a nucleon current correlator $\langle N(x)\,\bar{N}(0)\rangle$. This is especially useful for constructing gauge-invariant correlation functions: While the correlation functions with open quark or gluon legs depend on the gauge, hadronic current correlators are by construction gauge-invariant and thus in principle measurable.
@@
An important relation in QFT says that physical states produce poles in correlation functions (see e.g. Weinberg:1995mt for a proof). Suppose we split the $n$ coordinates in Eq. (4.6) into two sets $\{x_{i}\}$ and $\{x_{j}\}$ with $i=1\dots N$. Then, by inserting the completeness relation of the state space, each asymptotic one-particle state $|\lambda\rangle$ with onshell momentum $P$ and mass $m_{\lambda}$ (such that $P^{2}=m_{\lambda}^{2}$) produces a pole in any correlation function that is compatible with the quantum numbers of the state:
@@
which is the overlap of the vacuum and the one-particle state. Here, an overall coordinate $z$ has been factored out by translation invariance, so the BSWF depends only on N1N-1 coordinates. $\Psi$ is the QFT analogue of a quantum-mechanical wave function since it carries the information about the hadron, although it does not have a direct probability interpretation (more on that in Sec. 4.5). Note also that the BSWF is not an observable as it is gauge- and renormalization-point dependent.
@@
Figure 4.3: Left: Poles and cuts in the complex energy plane of a correlation function. By tuning certain control parameters, resonances may turn into bound states and vice versa. Right: Baryon poles appear in any correlation function that can produce them, such as the $N\pi$ scattering amplitude, the gauge-dependent qqqqqq correlation function or the respective gauge-invariant current correlator.
Is some given baryon resonance more like a three-quark state or a meson-baryon dominated state? A priori, QFT does not directly tell us that: From Fig. 4.3, a nucleon resonance can be thought of as an $N\pi$ state, but also as a three-quark state (qqqqqq), or a five-quark state ($qqqq\bar{q}$) etc., because all those correlation function must produce the same spectrum. For example, any dynamically generated ‘additional’ baryon resonance that arises from meson-baryon interactions in an EFT must also have a three-quark core (unless some quantum number prevents it), so these two options are not binary. Vice versa, a pole in the qqqqqq correlation function emerging from quark-gluon interactions must also appear in the $qqqq\bar{q}$ and other correlation functions, but the three-quark component might be small.
The relevant question is therefore: How much overlap does a state have with a given configuration? This is defined by the respective BSWF. If the qqqqqq overlap is small and the $qqqq\bar{q}$ overlap large, it implies a dominant $qqqq\bar{q}$ (or $N\pi$) structure; if the qqqqqq overlap is large it implies a three-quark dominance. In light-quark baryon spectroscopy, the $\Lambda(1405)$ is the primary candidate for having a dynamically generated additional state, which would then translate into two poles that couple more strongly to qqqqqq or $qqqq\bar{q}$ Meissner:2020khl , Mai:2020ltx , Mai:2022eur , Guo:2023wes . Going further, a deeper question is: What is the reason for the existence of a particular resonance? What are the interactions that produce a qqqqqq-like or molecular behavior? Is there a molecular force that binds $N$ and $\pi$, and this bound state may also couple to qqqqqq? Or is there a qqqqqq bound state that acquires a molecular component? Microscopic calculations that can extract such information using different approaches are underway and will presumably shed more light on these questions in the future. In addition, much can be learned from the structure of baryons as measured in electroproduction, which we will discuss in Sec. 5.
@@ 4.2  Effective field theories
Effective field theories (EFTs) for describing hadron-hadron interactions have found widespread applications over the last decades, see e.g. Guo:2017jvc , Oller:2019opk , Mai:2022eur , Doring:2025sgb for recent reviews. The primary quantities of interest are hadronic scattering amplitudes like in the top right row of Fig. 4.3. These contain the hadron poles matching the quantum numbers allowed by the initial and final states, which allows one to extract the hadron spectrum from the pole positions. Generally, the analytic structure of scattering amplitudes is constrained by the S-matrix principles of unitarity, analyticity and crossing symmetry. The analytic structure is determined entirely by physical particles: bound states and resonances manifest as poles, while particles in the loop diagrams generate branch cuts. By contrast, confinement implies that quarks and gluons must conspire in such a way that all direct evidence of their existence is erased from the singularity structure of physical scattering amplitudes. Although conceptually inevitable, the elimination of these gauge-dependent degrees of freedom from gauge-invariant amplitudes is quite astonishing and a central theme of quark-hadron duality Melnitchouk:2005zr .
@@
In practice the situation can be much more complicated than this simple sketch. For example, the case with two branch points is shown in Fig. 4.5. The four Riemann sheets in the left panel can be mapped to a torus Mai:2022eur in the right panel: The bottom and top borders of the rectangle are identified, and the left and right borders correspond to $|s|=\infty$. Crossing over between the four sheets is now an analytic operation.
@@
$\tau(s)=\frac{1}{16\pi}\sqrt{\frac{s-4m^{2}}{s}}=\frac{1}{16\pi}\frac{2|{\mbox{$q$}}|}{\sqrt{s}}\,$, (4.12)
[…]
@@
Note that the normalization conventions for $f_{l}$ and $K_{l}$ can differ in the literature.
@@
It is also often useful to perform a split $\mathbf{G}_{0}=\mathbf{G}_{1}+\mathbf{G}_{2}$ and define a quantity $\mathbf{K}_{1}^{-1}=\mathbf{K}^{-1}-\mathbf{G}_{1}$ $\Leftrightarrow$ $\mathbf{K}_{1}=\mathbf{K}+\mathbf{K}\,\mathbf{G}_{1}\,\mathbf{K}_{1}$, so that Eq. (4.15) turns into
@@
How is the kernel $\mathbf{K}$ determined, which contains the interactions between the particles and is the input in the equations? Since the particles appearing in the loops in Fig. (4.6) are hadrons, these interactions cannot be the ones between quarks and gluons but should rather be derived from an effective Lagrangian containing physical fields like nucleons, pions, etc. This leads to EFTs, whose prime example is chiral perturbation theory (ChPT) Gasser:1983yg , Gasser:1984gg , Weinberg:1978kz . In the two-flavor case, the ChPT Lagrangian is expressed in terms of pions which couple to nucleons; the extension to three flavors is straightforward. The idea behind ChPT is to be agnostic about the microscopic interactions and include all possible terms compatible with the symmetries of QCD – in particular chiral symmetry – which leads to a systematic expansion in powers of derivatives and pion masses. The elementary Feynman diagrams are shown in Fig. 4.7 together with the full correlation functions like the nucleon and pion propagators, $N\pi$ and $\pi\pi$ scattering amplitudes, etc., which can be calculated through perturbative loop expansions. The price to pay is that the Lagrangian is non-renormalizable and contains infinitely many terms with infinitely many free parameters. However, the derivative couplings ensure that higher loop diagrams in the perturbative expansion are suppressed at low momenta and small pion masses. The renormalization can then be done at each order and requires only a finite number of parameters at each order, the low-energy constants, which must be determined from elsewhere, e.g., from experiment or lattice QCD. In this way, ChPT is a low-energy EFT and one can make a range of predictions at low momenta and small pion masses, e.g. for the nucleon mass as a function of the pion mass, the $N\pi$ scattering amplitude, electroweak processes and so on, see Meissner:1993ah , Ecker:1994gg , Pich:1995bw , Bernard:1995dp , Scherer:2002tk , Bernard:2006gx , Bijnens:2006zp , Geng:2013xn for reviews.
@@
On the other hand, because ChPT amounts to perturbative loop expansions it cannot describe bound states and resonances. One strategy to deal with this is to include these extra fields ($\rho$ and $\omega$ mesons, the $\Delta$ resonance, Roper resonance, $\Lambda(1405)$, etc.) explicitly in the Lagrangian and hence in the perturbative expansions, see e.g. Meissner:1987ge , Ecker:1989yg , Jenkins:1991es , Hemmert:1997ye , Lutz:2001yb , Pascalutsa:2002pi , Bernard:2003xf , Kolomeitsev:2003kt , Pascalutsa:2006up , Djukanovic:2009zn , Gegelia:2016xcw and references therein. Another strategy is to unitarize ChPT by employing BSEs Oller:1997ng , Oller:1997ti , Oller:1998hw , Nieves:1999bx , Bruns:2010sv , Inoue:2001ip . A BSE automatically produces the unitary cuts when the hadrons in the loops go onshell, but because it is genuinely non-perturbative it can also dynamically generate bound states and resonances. This non-perturbative nature can be understood by analogy with the geometric series: The series expansion $f(x)=1+x+x^{2}+x^{3}+\dots$ determines the function f(x)=1/(1x)f(x)=1/(1-x) only for $|x|<1$, whereas the ‘nonperturbative’ equation
@@
where $f(x)$ also appears on the right, is always exact irrespective of how large $x$ is. Similarly, for small couplings one can expand Eq. (4.15) in $\mathbf{T}=\mathbf{K}+\mathbf{K}\,\mathbf{G}_{0}\,\mathbf{K}+\dots$, which does not generate any poles, while the original nonperturbative equation generates poles if the condition $\mathbf{K}\,\mathbf{G}_{0}=1$’ is satisfied for some value of $\sqrt{s}=m_{\lambda}$.
@@
A shortcoming of the BSE in the form of Eq. (4.15), which is only iterated in the $s$ channel, is that it does not automatically satisfy crossing symmetry. While it generates the correct $s$-channel cut for $s>4m^{2}$ (the unitary or right-hand cut), a unitarization also in the crossed channels would produce further cuts in the Mandelstam variables $t$ and $u$, which translates to $s<0$ and is the so-called left-hand cut. For this reason, other unitarization methods based on dispersion relations such as the $N/D$ approach Chew:1960iv , Oller:1998zr , Meissner:1999vr , the inverse amplitude method Truong:1988zp , Pelaez:2006nj , GomezNicola:2007qj , Pelaez:2010fj or the Roy-Steiner equations Hoferichter:2015hva , Hoferichter:2015tha , RuizdeElvira:2017stg , Hoferichter:2023mgy have been frequently employed, see e.g. the reviews Pelaez:2010fj , Hoferichter:2015hva , Mai:2022eur for details and references.
@@
The application of BSEs with kernels taken from EFTs, with (multi-)pion and possibly other hadron exchanges, has found widespread applications in hadron and nuclear physics. Although we exemplified it for the two-body scattering amplitude in a single channel, the approach can be generalized to dynamical coupled-channel equations where the external particles create different intermediate states such as $N\pi$, $\Lambda K$, $\Sigma K$, $N\gamma$, etc., see Doring:2025sgb for a recent review. Coupled-channel models have become indispensable tools for amplitude analyses to extract experimental data, especially in view of extracting the photo- and electrocouplings of light baryons (see also the discussion in Sec. 2.2). Examples are the ANL-Osaka approach Sato:1996gk , Matsuyama:2006rp , Julia-Diaz:2007qtz , Suzuki:2009nj , Kamano:2013iva , the Jülich-Bonn-Washington approach Doring:2010ap , Ronchen:2012eg , Ronchen:2014cna , Wang:2022osj , Mai:2021vsw , Wang:2024byt , Doring:2025sgb , and various K-matrix formulations like the MAID Tiator:2011pw , Drechsel:2007if , Drechsel:1998hk , SAID Arndt:2002xv , Workman:2012jf , JLab Aznauryan:2011qj , KSU Manley:1992yb , Shrestha:2012ep , Giessen Penner:2002ma , Penner:2002md , Shklyar:2006xw , Shklyar:2012js and Bonn-Gatchina approaches Anisovich:2005tf , Anisovich:2010mks , Anisovich:2011fc , Anisovich:2017bsk , Sarantsev:2025lik .
@@
Dynamical coupled-channel models also allow one to estimate the meson-baryon dressing effects of nucleon resonances and extract their pole trajectories in the complex plane, see e.g. Suzuki:2009nj , Guo:2023wes . As an example, the procedure employed in the ANL-Osaka approach is illustrated in Fig. 4.8 for a single-channel $N\pi$ scattering equation. If one pulls out the resonant $s$-channel $N^{\ast}$ contributions from the kernel explicitly, then also the T-matrix turns into the sum of a non-resonant and a resonant part. The former satisfies the scattering equation with the non-resonant kernel, while the dressing effects for the resonance coupling and propagator are calculated separately. In this way one can study the effects of meson-baryon interactions: When the full T-matrix is fitted to the electroproduction data, switching off the dressing effects for the masses and transition form factors provides an estimate for the ‘quark core’ of the resonance.